Rochester Heating Systems, Inc. calculates its predetermined overhead rate on a quarterly basis. The following estimates were made for the current year. Quarterly Predetermined Overhead Estimated Estimated Rate (per Manufacturing Direct-Labor direct-labor Hours 25,000 Overhead hour) First quarter 75,000 Second 16,000 12,500 14,000 96,000 quarter Third quarter 37,500 Fourth 84,000 quarter Total $292,500 67,500 The firms main product, part number SC71, requires $200 of direct material and 20 hours of direct labor per unit. The labor rate is $15 per hour2. Determine the cost of one unit of part number SC71 if it is manufactured in February versus May. February May Total cost per unit
Answer #1

Answer

  • All working forms part of the answer
  • Workings

Estimated Manufacturing Overhead

Estimated Direct Labor Hours

Predetermined Overhead rate per DLHs

[A]

[B]

[C = A/B]

Quarter 1 [Jan, Feb, Mar]

$75,000

                  25,000

$3 [to be used for February]

Quarter 2 [April, May, June]

$96,000

                  16,000

$6 [to be used for May]

Quarter 3

$37,500

                  12,500

$3

Quarter 4

$84,000

                  14,000

$6

  • Requirement 2

--February [FALLS UNDER QUARTER #1]

Total Cost = $ 200 material + (20 hours x $15 per hour) labor + (20 hours x $ 3 per DLH) overhead
= 200 + 300 + 60
= $ 560 = Answer

--May [FALLS UNDER QUARTER #2]

Total Cost = $ 200 material + (20 hours x $15 per hour) labor + (20 hours x $ 6 per DLH) overhead
= 200 + 300 + 120
= $ 620 = Answer